# High-Voltage Performance Testing in LAr of the PMMA Cathode Connection for the DarkSide-20k Experiment

### v1

@@ 2 Electric Fields and High–Voltage in DS–20k
In DS–20k, high voltage is delivered to the cathode using a custom HV cable rated up to kV and manufactured by Dielectric Sciences (Lowell, MA, USA). The cable features a three–layer polyethylene (PE) structure, with a carbon–loaded central conductor (2 mm diameter) surrounded by a mm thick PE insulating layer and a mm thick outer carbon–loaded PE ground. The HV cable enters the DS–20k vessel and runs along the full height of the TPC, connecting to the underside of the cathode via a dedicated stress–cone assembly (Fig. 1(a)), consisting a PE–cone plug, and a PMMA–cone bonded to the cathode PMMA window.
The PE–cone consists of an insulating PE sleeve and a carbon–loaded conductive PE outer shell, thermally fused onto the cable and fabricated by the UC Davis group. The cable terminates at a copper connector mounted on the inner surface of the cathode, in direct electrical contact with the Clevios conductive coating.
The highest electric–field regions are located at the PE–cone/PMMA–cone interface and beneath the copper connector. A COMSOL study (Fig. 1(b)) with the cathode biased at kV shows that the electric field in these regions remains below the breakdown threshold in LAr ( kV/cm [3]), corresponding to a safety margin of approximately with respect to the nominal cathode voltage. Under these conditions, the cathode HV system must operate while minimizing the risk of electrical discharges, ensuring a reliable connection within a constrained geometry, and maintaining stability at liquid–argon temperature ( K).
@@ 3 Experimental Test Setup at UC Davis
A COMSOL simulation was used to validate the consistency of the test setup with the DS–20k cathode plug in terms of local electric–field conditions (Fig. 2(b)). With the aluminum sphere biased at kV, the electric fields in the critical regions were found to be comparable to those expected in DS–20k.

### v2

@@ 2 Electric Fields and High–Voltage in DS–20k
In DS–20k, high voltage is delivered to the cathode using a custom HV cable rated up to kV and manufactured by Dielectric Sciences (Lowell, MA, USA). The cable features a three–layer polyethylene (PE) structure, with a carbon–loaded central conductor (2 mm diameter) surrounded by a mm thick PE insulating layer and a mm thick outer carbon–loaded PE ground. The three PE layers are co-extruded during the cable manufacturing process. The HV cable enters the DS–20k vessel and runs along the full height of the TPC, connecting to the underside of the cathode via a dedicated stress–cone assembly (Fig. 1(a)), consisting a PE–cone plug, and a PMMA–cone bonded to the cathode PMMA window.
The PE–cone consists of an insulating PE sleeve and a carbon–loaded conductive PE outer shell, thermally fused onto the cable and fabricated by the UC Davis group. Along the cable section extending inside the stress cone toward the cathode connection, the outer ground layer is mechanically stripped by longitudinal cutting and peeling to expose the insulating PE layer. The cable terminates at a copper connector mounted on the inner surface of the cathode, in direct electrical contact with the Clevios conductive coating.
The highest electric–field regions are located at the PE–cone/PMMA–cone interface and beneath the copper connector. A COMSOL study (Fig. 1(b)) with the cathode biased at kV shows that the electric field in these regions remains below the breakdown threshold in LAr ( kV/cm [3]), corresponding to a safety margin of approximately with respect to the nominal cathode voltage. The outer surface of the PMMA cylinder is coated with Clevios and held at ground potential. The simulation does not include spatial or time–dependent surface charge accumulation effects. Under these conditions, the cathode HV system must operate while minimizing the risk of electrical discharges, ensuring a reliable connection within a constrained geometry, and maintaining stability at liquid–argon temperature ( K).
@@ 3 Experimental Test Setup at UC Davis
A COMSOL simulation was used to validate the consistency of the test setup with the DS–20k cathode plug in terms of local electric–field conditions (Fig. 2(b)). With the aluminum sphere biased at kV, the electric fields in the critical regions were found to be comparable to those expected in DS–20k. Surface charge accumulation effects are not included in the simulation.